### v1

@@ 4.3 The positivity preserving interpolant
$\mathbf{I}_{1}$ $\leq\int_{Q}\omega(x)|{\mathbf{w}}(x)-{\mathbf{w}}(\hat{x})|^{2}\,\mbox{\rm d}x=\int_{Q}\omega(x)\left|\int_{0}^{1}\nabla{\mathbf{w}}(tx+(1-t)\hat{x})\cdot(x-\hat{x})\right|^{2}\,\mbox{\rm d}t\,\mbox{\rm d}x$
@@
where we have used the fact that $\omega$ belongs to the reduced class $A_{1}(\Omega)$ (see Definition 2.5). We now introduce a coordinate system in which the first axis is aligned with $\boldsymbol{\zeta}$ and $Q$ is on the half space defined by $\{x\in\mathbb{R}^{d}:x_{1}\geq 0\}$. We thus have that, if $x\in Q$, then $x=(x_{1},x^{\prime})^{\top}$ with $x_{1}\geq 0$ and $x^{\prime}\in F\subset\mathbb{R}^{d-1}$. We also note that $\hat{x}=(0,x^{\prime})^{\top}$, $x-\hat{x}=(x_{1},0)^{\top}$, and $y=tx+(1-t)\hat{x}=(tx_{1},x^{\prime})^{\top}$ (see Figure 1). With these observations in mind, we continue the bound for the term $\mathbf{I}_{1}$ as follows:
@@
where the hidden constant depends on $[\omega]_{A_{1}}$ and $\|\omega\|_{L^{\infty}({\mathcal{G}})}$.
@@
Figure 1: The geometric setting for the bound for $\mathbf{I}_{1}$ in the proof of Proposition 4.7. The element $T\in\mathscr{T}_{h}^{\partial}$ is shown in black, while the related prism is in dashed blue. We also show a generic point $x\in Q$ and its projection onto $F$, i.e., $\hat{x}\in F$.
4.4 Error estimate
We are now ready to present an error estimate for the obstacle problem.
@@ 4.4 Error estimate
Let $u\in W^{1,\varphi}_{0}(\omega,\Omega)$ and $u_{h}\in\mathbb{V}_{h}$ solve (39) and (44), respectively. Assume that $u,{\mathbf{V}}(\nabla u)\in W^{1,2}(\omega,\Omega)$, that $\lambda$, defined in (40), belongs to $L^{2}(\omega^{-1},\Omega)$, and that $\omega\in A_{1}(\Omega)$. If $\mathscr{T}_{h}$ is such that no simplex has more than one $(d-1)$-dimensional subsimplex on $\partial\Omega$ and $h$ is sufficiently small, then
$\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}\lesssim h\|\nabla{\mathbf{V}}(\nabla u)\|_{L^{2}(\omega,\Omega)}+h^{\frac{1}{2}}\|\nabla u\|_{L^{2}(\omega,\Omega)},$
@@
Proof 4.10.
We apply Proposition 4.7 to the conclusion of Proposition 4.3 and obtain
@@
It remains to invoke the approximation property of the interpolant $\wp_{h}$, which was proved in Lemma 4.5, in order to be able to conclude.
@@
Note that the error estimate depends on the assumption that ${\mathbf{V}}(\nabla u)\in W^{1,2}(\omega,\Omega)$. In addition, we need the regularity assumption $\lambda\in L^{2}(\omega^{-1},\Omega)$. We leave the exploration of these properties as open conjectures and refer the reader to [9, 10, 20, 25, 29, 45] for some results in this direction. We also note that the assumption $\nabla u\in L^{2}(\omega,\Omega)$ is not extremely restrictive. In the case of the $p$-Laplacian, for example, this reduces to the fact that $p\geq 2$. Indeed,
$\int_{\Omega}|\nabla u(x)|^{2}\omega(x)\,\mbox{\rm d}x\leq\left(\int_{\Omega}|\nabla u(x)|^{p}\omega(x)\,\mbox{\rm d}x\right)^{2/p}\left(\int_{\Omega}\omega(x)\,\mbox{\rm d}x\right)^{\frac{p-2}{p}}.$

### v2

@@ 5.3 The positivity preserving interpolant
$\mathbf{I}_{1}$ $\leq\int_{Q}\omega(x)|{\mathbf{w}}(x)-{\mathbf{w}}(\hat{x})|^{2}\,\mbox{\rm d}x=\int_{Q}\omega(x)\left|\int_{0}^{1}\nabla{\mathbf{w}}(tx+(1-t)\hat{x})\cdot(x-\hat{x})\,\mbox{\rm d}t\right|^{2}\,\mbox{\rm d}x$
@@
where we have used the fact that $\omega$ belongs to the reduced class $A_{i(\Phi)}(\Omega)$ (see Definition 2.1). We now introduce a coordinate system in which the first axis is aligned with $\boldsymbol{\zeta}$ and $Q$ is on the half space defined by $\{x\in\mathbb{R}^{d}:x_{1}\geq 0\}$. We thus have that, if $x\in Q$, then $x=(x_{1},x^{\prime})^{\top}$ with $x_{1}\geq 0$ and $x^{\prime}\in F\subset\mathbb{R}^{d-1}$. We also note that $\hat{x}=(0,x^{\prime})^{\top}$, $x-\hat{x}=(x_{1},0)^{\top}$, and $y=tx+(1-t)\hat{x}=(tx_{1},x^{\prime})^{\top}$ (see Figure 1). With these observations in mind, we continue the bound for the term $\mathbf{I}_{1}$ as follows:
@@
where the hidden constant depends on $\omega]_{A_{i(\Phi)}}$ and $\|\omega\|_{L^{\infty}({\mathcal{G}})}$.
@@
Figure 1: The geometric setting for the bound for $\mathbf{I}_{1}$ in the proof of Proposition 5.7. The element $T\in\mathscr{T}_{h}^{\partial}$ is shown in solid black, while the related prism is in dashed blue. We also show a generic point $x\in Q$ and its projection onto $F$, i.e., $\hat{x}\in F$.
5.4 An error estimate
We conclude with the following error estimate for the obstacle problem. Recall that we have assumed that the obstacle satisfies $\psi\in C^{2}(\bar{\Omega})$.
@@ 5.4 An error estimate
Let $\Phi$ be an $N$-function with $i(\Phi)>2$ that satisfies Assumption 3, and let $\omega\in A_{2}(\Omega)$. Let $u\in W^{1,\Phi}_{0}(\omega,\Omega)$ and $u_{h}\in\mathbb{V}_{h}$ solve (46) and (51), respectively. Assume that ${\mathbf{V}}(\nabla u)\in W^{1,2}(\omega,\Omega)$, and that $\lambda$, defined in (47), belongs to $L^{2}(\omega^{-1},\Omega)$. If $\mathscr{T}_{h}$ is such that no simplex has more than one (d1)(d-1)-dimensional subsimplex on $\partial\Omega$ and $h$ is sufficiently small, then
$\|{\mathbf{V}}(\nabla u)-{\mathbf{V}}(\nabla u_{h})\|_{L^{2}(\omega,\Omega)}\lesssim h\|\nabla{\mathbf{V}}(\nabla u)\|_{L^{2}(\omega,\Omega)}+h^{\frac{1}{2}}\|\nabla(u-\psi)\|_{L^{2}(\omega,\Omega)},$
@@
Proof 5.10.
We begin with the observation that, since $i(\Phi)>2$, we have that $A_{2}\subset A_{i(\Phi)}$, and thus that $\omega\in A_{i(\Phi)}(\Omega)$. We may thus apply Proposition 5.7 to the conclusion of Proposition 5.3 and obtain
@@
It remains to use the approximation property of the positivity preserving interpolant $\wp_{h}$ proved in Lemma 5.5 to be able to conclude.
@@
Note that the error estimate depends on the assumption that ${\mathbf{V}}(\nabla u)\in W^{1,2}(\omega,\Omega)$. In addition, we need the regularity assumption $\lambda\in L^{2}(\omega^{-1},\Omega)$. We leave the exploration of these properties as open conjectures and refer the reader to [9, 10, 20, 25, 28, 48] for some results in this direction.